OpenStudy (butterflyhope):

Urgent *will medal*: Choose a variable to represent the unknown in this problem. Write an equation and solve: Mr. James is looking for the best deal on tickets for the concert. At the box office he would have to pay $15.75 for a ticket. On the Internet he could purchase tickets for 4/5 of that price. How much would 6 tickets cost over the Internet? Thanks c:

OpenStudy (anonymous):

We don't know the cost so we will label that "c" we want to know how much 6 tickets will cost for 4/5 the price of tickets at 15.75 right?

OpenStudy (butterflyhope):

Yeah... Hm...would it be written as an inequality?

OpenStudy (anonymous):

That could work. but i feel like we could make it easier by trying to simplify what we already have

OpenStudy (anonymous):

the price of a ticket online is 4/5 that of 15.75 what would be an easier way to write that?

OpenStudy (anonymous):

Let the cost of the ticket on the internet be x. For 1 ticket on the internet, 4/5 ($15.75) = $12.60. You need 6 tickets, Therefore 6 ($12.60) Answer is $ 75.60

OpenStudy (anonymous):

We could say that a ticket online is 4/5 (15.75)= 12.60 We have the price for a ticket online and want to know the total cost of multiple tickets along with c, we could use t for tickets and use the equation y=ct we know c which is 12.60 y=12.60t t=6 y=(12.60)(6) y=$75.60